Thursday 06/21/12

2:00P — 8:30P

  • Calibrating new bead imaging: 1:20 then 1:500 540 beads + 1:500 Q-dots-705.  Too dense.
  • Dilute a further 1:500.  Q dots good density.  Blinky, excited best by 405 or 488.  Look good in 647, not really visible in 750 channel — more visible in 561 than 750.   🙁
  • Hard to unmix the 488 excited Q-dots-705 from the 540 beads (potentially feasible by intensity).
  • 3D bead measurement not mapping IR 647 and Vis 647 to the same plane.
  • mites detected in fly room!  Evacuated fly stocks to Natio for time being.  Keep under observation!
  • STORM imaging of Antp locus and H3-750 with weak signal from H3K27-647.

 

This entry was posted in Summaries and tagged , , . Bookmark the permalink.